### v2

# LAP, Using Action Feasibility for Improved Uncertainty Alignment of Large Language Model Planners

Abstract
Large language models (LLMs) showcase many desirable traits for intelligent and helpful robots. However, they are also known to hallucinate predictions. This issue is exacerbated in robotics where LLM hallucinations may result in robots confidently executing plans that are contrary to user goals, relying more frequently on human assistance, or preventing the robot from asking for help at all. In this work, we present LAP, a novel approach for utilizing off-the-shelf LLMs, alongside a novel Action feasibility metric, in robotic Planners that minimize harmful hallucinations and human intervention. Our key finding is that calculating and leveraging a new metric, which we call A-Feasibility, a measure of whether a given action is possible and safe in the provided scene, helps to mitigate hallucinations in LLM predictions and better align the LLM’s confidence measure with the probability of success. We specifically propose an A-Feasibility metric which both combines scene context and prompting a LLM to determine if a given action is possible and safe in the scene, using the LLM’s response to compute the score. Through experiments in both simulation and the real world on tasks with a variety of ambiguities, we show that LAP significantly increases success rate and decreases the amount of human intervention required relative to prior art. For example, in our real-world testing paradigm, LAP decreases the human help rate of previous methods by over 33% at a success rate of 70%.
I Introduction
Figure 1: LLMs are not grounded in the real world and will frequently hallucinate. Additionally, when provided with an ambiguous instruction, the LLM must know when to ask users for help. We present LAP, a novel approach for calibrating LLM confidence with an action feasibility metric to better detect hallucinations and resolve ambiguities, before asking a user for help when necessary. We calculate a novel metric, which we call A-Feasibility, for each possible action and combine it with the next token probability of that action to determine both which option is most likely, and if help is needed. Note, in the provided example, ‘pick up coffee cup’ is a hallucination caused by the users phrasing of their command. No such specific object was grounded in the provided ground truth perception information. In LAP, our A-Feasibility metric, seen in orange, mitigate this case.
Imagine you have a home robot and you want it to bring you your coffee cup. As you tell the robot your instruction, it should comprehend your goal, no matter how you phrase the instruction, and ideally complete the task without further clarification. Now imagine there are multiple coffee cups on the counter, the robot may need to ask you for help determining which one is yours. Finally, imagine that there is one coffee cup, but many other objects on the counter that may distract the robot. You would expect the robot to get the coffee cup, without asking for your assistance. If the robot asked you for help, brought you the wrong item, or failed due to some hallucination, you would be inclined to question its ability and potentially be less likely to trust it in the future. There will always be uncertainty in the unstructured and novel environments these types of robots operate in, but they must operate reliably and intelligently nonetheless.
Recent approaches that leverage large language models (LLMs) for planning [1, 2, 3] have demonstrated an ability to navigate these types of complex environments with a higher success rate than prior methods, while also responding properly to natural and unstructured language instructions. Furthermore, each new generation of the language model, such as GPT-3 to GPT-4, greatly improves model capabilities and intelligence, and subsequently performance on these robotics tasks. However, a significant challenge with all LLMs, new or old, is their tendency to present an incorrect answer confidently, or hallucinate [4, 5].
Additional challenges arise when the LLM is embodied and must interact with users directly. Significant uncertainty is present as human-provided instructions can be ambiguous, in some cases causing further hallucinations. Completing the wrong action or hallucinating and failing could hurt user trust or even be dangerous. One method of mitigating these issues is to ask users for help or clarifications when needed [3, 6, 7, 8]. However, most prior work does not ask for clarification from users in uncertain or ambiguous situations, or does so via extensive manual programming, oftentimes excessively relying on seeking assistance [6]. More recently, KnowNo [3] frames the task of when a robot should ask for help as uncertainty alignment, and outlines the main challenges for robots that ask for help: (i) calibrated confidence – the robot should seek sufficient help to ensure a probability of task success, and (ii) minimal help – the robot should minimize the amount of help it seeks. However, KnowNo primarily focuses on creating a means of evaluation for the uncertainty alignment task and provides a basic baseline. In contrast, we aim to improve the confidence measure, increase the success rate and minimize human intervention, moving towards more helpful and less annoying robot partners.
Main Contributions: We introduce LAP, a novel approach for calibrating LLM confidence with action feasibility. Our key finding is that a new metric, which we call A-Feasibility, which can intuitively be thought of as the potential for an action to be both possible and safe in the given scene, can minimize hallucinations, improve safety, and more accurately calibrate uncertainty. By handling these issues, the A-Feasibility metric can be used to increase the task success rate and minimize the frequency of human intervention. An additional advantage of our A-Feasibility metric is that it does not require extensive training, which is difficult to do without overfitting for extremely diverse and data-limited tasks like those found in home robotics.
The main contributions of our work include:
  1. 1.
    We introduce LAP, a novel approach for uncertainty alignment that uses Large Language Models and a feasibility to better align model confidence with task success for high-level robotic planners.
  2. 2.
    We evaluate our approach in both simulation and on hardware in the real world using a suite of language-instructed manipulation tasks from the KnowNo Simulation and Mobile Manipulator Datasets [3]. We show that LAP significantly increases the success rate and reduces the amount of help needed as compared to our baselines across different environments and LLMs. Specifically in our real-world testing, we see a decrease in the human help rate of previous methods by over 33% at a success rate of 70%, with a similar decrease at most success rates.
  3. 3.
    We explore and evaluate different ways of extracting the A-Feasibility metric for use in LAP. We propose the use of a combined metric which utilizes scene context alongside a novel prompt-based metric, which allows added flexibility in the A-Feasibility metric to encapsulate safety and other important behavior.
  4. 4.
    We show that using LAP on the newest, most powerful language models out-of-the-box outperforms fine-tuning LAP and prior art [3] on the best models available for fine-tuning on this task.
Figure 2: We present LAP, a novel approach for uncertainty alignment that utilizes a novel A-Feasibility metric. After generating a set of multiple-choice candidate actions, each prepended with ‘A’, ‘B’, ‘C’, etc., we calculate the A-Feasibility metric for each option (shown in orange). One version of our method uses the scene context from earlier prompts (shown in blue), while another sends a new query to the LLM to score a given option. This prompt-based method allows us to include factors like safety into the A-Feasibility metric. After computing each option’s A-Feasibility metric, we multiply the next-token likelihood and construct a prediction set of options with a total value above a threshold, $t$. Note that we eliminate option A, ”Put metal bowl in the microwave,” as a valid option as it is dangerous. If the size of the prediction set were greater than one, the agent would ask for help.
II Related Work
II-A Hallucinations and Uncertainty in LLMs
As LLMs have risen to the fore, an increasing amount of work has been aimed at addressing their hallucinations, times when the LLM generates content not grounded in reality. Some works aim to quantify the frequency or severity of hallucinations [4] in a given LLM, while many aim to decrease their frequency [5], or better calibrate uncertainty [9, 10, 11]. Some recent works attempt to handle hallucinations by using conformal prediction-based methods to provide coverage guarantees[12]. Ren et al. [3] introduce KnowNo, which uses conformal prediction to create coverage guarantees for robotics tasks, asking users for help when the agent is unsure. In contrast to these works, we mitigate hallucinations through the introduction of our A-Feasibility metric, minimizing the amount of help needed.
II-B Feasibility and Robotics
Affordances, or what is possible given environment and object characteristics, is a well-studied topic in computer vision and graphics [13, 14, 15]. Most of these works focus on placing virtual humans or animations into 3D scenes such that the resultant placement generates visually plausible results [14, 16, 17]. Recent works in robotics use similar techniques to determine what actions are possible and likely for a robot agent given the environmental context, typically for manipulators[18]. SayCan [1] applied these techniques to robot planners by adding an affordance value, obtained from trained value functions and indicating the value of an action in the environment, into the planning scheme.
We take inspiration from the use of affordances and extend its definition to what we call action feasibility, or the potential to complete an action in the given scene safely and effectively. We also computer our A-Feasibility metric without the need for training or LLM fine-tuning. We choose to do this as training capable value functions for every possible action in the real world (as SayCan does) is extremely difficult due to a lack of domain data and vast task diversity. We also show that using feasibility can help mitigate hallucinations in an LLM’s generations by grounding the generations in the real world.
II-C LLMs for Robot Planning and Human-Robot Interaction
Robot motion planning is a well-studied problem in robotics [19, 20, 21]. Recently, Large Language Models (LLMs) have shown an ever-increasing set of capabilities from reasoning and logic [22, 23, 24] to math and physics [25]. This has extended to robotics, with LLMs improving the state of the art in tasks from high-level planning [1, 26, 27, 28], to object goal navigation[29, 2].
Natural language is one of the most popular methods of producing effective human-robot interaction [30, 7, 31, 8] and the advent of LLMs has made dialogue an even more natural medium of communication between a user and their robotic agent. Recent LLM-based robotic works have extended their tasks, incorporating user interaction [6, 32, 3] to receive user instructions or to resolve uncertainty.
III A-Feasibility Metric for Decreasing Uncertainty
In this work, we consider the following problem statement: Given a robotic agent and a user-provided instruction, execute the instruction completely and accurately while resolving any ambiguities as needed through dialogue with the user. Minimize the amount of dialogue. Specifically, our robot will begin in a set environment, awaiting a user’s instruction. The user will provide an instruction, which may be under-specified, at which point our method will need to evaluate whether or not it requires additional information from the user. This problem space was initially explored by [3]. Following [3], we choose to transform the problem into an multiple choice question answer (MCQA) task, which allows us to define the problem as next-token prediction – aligning well with LLM loss functions and training data. To do this, an LLM is provided a few-shot prompt (provided in supplementary materials) that includes possible next steps in a few example scenarios. The LLM in turn generates a set, $\{y^{i}\}$, of candidate actions that are semantically different for the current scenario. Then, the task of choosing among the candidate actions is formatted as a multiple choice question to the LLM, with each option preceded with ‘A’, ‘B’, ‘C’, ‘D’, or ‘E’ as shown in Figure 2. The probabilities of each possible token (i.e. the multiple choice options ‘A’, ‘B’, ‘C’, …) then serve as normalized scores that can be utilized by uncertainty quantification methods, like LAP and KnowNo [3]. We utilize these normalized scores alongside our A-Feasibility metric to generate a prediction set of plans from $\{y^{i}\}$. Please see [3] for further rationale on why MCQA should be used for uncertain robot planning.
Method Overview. We present an overview of our method, LAP, in Figure 2. LAP begins with the MCQA paradigm described above, outputting the model’s uncalibrated confidence score for each generated action option, $p^{*}(y^{i})$. We then compute the A-Feasibility metric, $a^{i}$, for each option, corresponding to the plausibility of that action given the scene. We describe three different methods of obtaining this score below, context-based, perception-based, and prompt-based. Each of these scores can be used independently or combined into a superscore as shown in Figure 2. We choose to present all three as each could be advantageous in different use cases and their combination can create a more performative final score. We now define our final probability of each option being the user’s intent as:
$p(y^{i})=p^{*}(y^{i})\cdot a^{i}.$ (1)
With each option scored, we generate the prediction set of plans, $\mathcal{P}$ as plans with a score above a threshold, $t$
$y^{i}\in\mathcal{P}\text{ if }p(y^{i})>t.$ (2)
The LLM is certain if this set is a singleton, and queries the user for help if not. Calculating an optimal $t$ is the main subject of KnowNo [3], which uses conformal prediction to provide statistical guarantees of a certain success rate. Our work differs significantly from KnowNo as we improve $p(y^{i})$ by integrating our A-Feasibility metric.
Context-Based A-Feasibility Metric. Our first method of calculating the A-Feasibility metric that we explore is the most efficient, requiring the least amount of time, compute, and control over the robot agent. This method uses the context of the scene that the robot already found and provided to the LLM from its perception algorithms. We label this A-Feasibility metric $a^{i}_{c}$. While perception information is provided, this does not always prevent the LLM from hallucinating new objects into existence.These types of hallucinations become even more prominent when the hallucinated object is in the user command. An example of this can be seen in Figure 1 with the ‘coffee mug’ object.
To calculate $a^{i}_{c}$, the scene context is first parsed into a set of found objects, $\mathcal{O}$. Each generated action candidate from the LLM’s MCQA creation response, $y^{i}$, is parsed for any objects, $o^{i}$, present. If all of the objects in $y^{i}$ are in the set of found objects, $a^{i}_{c}=1$. Otherwise, $a^{i}_{c}=0$. To formalize this:
$a^{i}_{c}=\begin{cases}1&\text{if }\forall o^{i}\in y^{i},\,o^{i}\in\mathcal{O}\\ 0&\text{otherwise}.\end{cases}$ (3)
Perception-Based A-Feasibility Metric. Building off of the context-based A-Feasibility metric, the perception-based A-Feasibility metric takes into account any uncertainty in the perception algorithms themselves. To do this, we run an open-vocabulary vision-language model, like CLIP [33] or ViLD [34], on the scene observations with each object, $o^{i}$ (from the option generated by the LLM, $y^{i}$) as input.
Specifically, we take the camera images of the scene, $s$, and feed them through ViLD with the list of objects $\{o^{i}\}$ as input. We then take the probabilities of each object $p(o^{i}|s)$ as outputted by the model and find the average for a given option, $y^{i}$. Formally, we define $a^{i}_{perc}$ as
$a^{i}_{perc}=\frac{1}{n}\sum_{i=1}^{n}p(o^{i}|s),$ (4)
for $n$ objects in $y^{i}$. If the object is localized to the same location as an object in the scene context using IoU (usually an indicator of a hallucination), the metric is lowered to 0 to prevent duplicate options of the same object.
Prompt-Based A-Feasibility Metric. Recent works like [5] and [22] have shown that additional prompts to the LLM requiring it to reason on its prior outputs can isolate harmful hallucinations. We leverage this idea and propose a prompt-based A-Feasibility metric for each action option, $y^{i}$, generated by the model. This additionally allows us to generate a more flexible A-Feasibility metric, which can account for factors beyond just the grounding of objects in the scene. Our prompt, given our real-world evaluation environment, begins with a series of few-shot samples, before serving the following text:
{Header}
We: On the counter, there is
$\mathcal{O}$}.
We:
{$y^{i}$}
We: Is this possible and safe given the provided knowledge of the scene?
You:
In the prompt, the items inside the brace notation are replaced with their respective values. The header includes the description of items found in the scene from the MCQA generation and scoring prompts and describes the task for the LLM: determining whether the given action is ‘possible and safe’ in the current scene. The header also provides a few examples. Once we receive the response from the LLM, we extract the normalized probability of the ‘True’ token and utilize this as our A-Feasibility metric $a^{i}_{pr}$, for the option $y^{i}$.
Figure 3: Comparison of task success rate vs human help rate (left) and average prediction set size (right) in our Simulation experiments. Note LAP improving upon KnowNo for almost all success rates. Including a reduction in the human help rate of over 70% at a success rate of 60%. No Help and Binary do not create prediction sets and are not included in the plot.
One benefit of this approach is the ability to tailor the A-Feasibility metric to new criteria. For example, in this prompt, we add safety as a major component of the metric, allowing the LLM to lower the metric for putting the metal bowl in the microwave. Similarly, if an important characteristic of the robot agent is that it does not leave a certain area, one could specify this and the A-Feasibility metric should adapt. The perception-based methods above cannot apply logic to the generated options, $y^{i}$, in this way.
Fine-Tuning. We also hypothesize that utilizing the newest, most powerful, LLM available off-the-shelf will generally outperform fine-tuning the older models available for fine-tuning. We tested this hypothesis by fine-tuning gpt-3.5-turbo-1106 [35], the most powerful model available from OpenAI for fine-tuning. For the simulation environment, we fine-tuned on each of the three LLM phases of LAP, the MCQA generation, MCQA scoring, and A-Feasibility scoring.
IV Experiments and Results
We evaluate LAP in a diverse set of language-instructed tasks and environments as described below. Each of these environments is outlined in [3]. We chose to use these environments for easy comparison to [3] and to continue to move towards a standard set of evaluations for the uncertainty alignment task. Each environment has a parameterized scenario distribution with a pre-defined set of possible ambiguities in the human’s instruction. We utilize the truth labels that they provide. The full outline of the dataset distribution and every possible ambiguity is included in the supplementary material. We utilize GPT-4 [36] as the LLM in all experiments and ablate against it separately.
Baselines. Our primary baseline is KnowNo [3], which uses the probability of each option directly from the LLM, and builds a prediction set of options with a probability greater than a threshold value. KnowNo uses conformal prediction to set the threshold value in such a way that a statistical guarantee of a certain success rate is possible. They do nothing to improve the confidence measure itself outside of the task framing as MCQA. All of the additional baselines use this MCQA framing. Other baselines include:
  • Prompt: Prompts the LLM to output the prediction set (e.g. “Prediction set: [B, D]”) Exact prompts are in the supplemental materials.
  • Binary: Prompts the LLM to directly output a binary indicator of uncertainty (e.g. “Certain/Uncertain: Certain”) similar to [6].
  • No Help: Always uses the highest score directly from the LLM without creating a prediction set or asking for human intervention.
Evaluation Metrics. We primarily use human-help rate and prediction set size versus success rate to compare LAP to baselines. Intuitively, when given a set success rate, these metrics provide information about the help rate and average prediction set size for that success rate, with lower help rates and lower average prediction set sizes better. For a given data sample, the human-help rate is 1 if the prediction set size is greater than 1. Prompt and Binary produce a single success rate and thus a singular help rate and average prediction set size. With KnowNo and LAP, we have a tunable threshold value that is used to confine the prediction set. For these methods, we apply threshold values, $t$, ranging from 0.0000001 to 0.7 for full coverage. The success rate and human help rate or average prediction set size for each threshold value are then plotted, and the area under the resultant curve (AuC) becomes a valuable metric to quantify an improved relationship between task success rate and human help rate.
Figure 4: Comparison of task success rate vs human help rate (left) and prediction set size (right) in our real-world experiments on a ClearPath TurtleBot. Note the improvement over KnowNo across all success rates. LAP with the prompt and context-based A-Feasibility metrics outperforms using each score separately, while all improve upon KnowNo.
IV-A Simulation: Tabletop Rearrangement
In this task, a robot arm is asked to rearrange objects on a table in the PyBullet simulator. Each scenario is initialized with three bowls and three blocks, with one each of blue, green, and yellow, respectively. The task requires the robot to move a certain number of blocks or bowls toward a specific location relative to a different object. For example, “Move the Green Block to the left of the Blue Bowl.” As introduced KnowNow, three different ambiguities are injected into the user instruction: (1) attribute, where ‘block’ or ‘bowl’ is replaced with a potentially ambiguous term like ‘thing’ or ‘cube,’ (2) numeric, where the number of blocks to move is underspecified by using a term like ‘some’ in place of ‘two,’ and (3) spatial, where the exact location like ‘left’ or ‘ front’ is replaced with an ambiguous term like ‘near.’
We show in Figure 3 that LAP with the context-based A-Feasibility metric ($a_{c}$) outperforms KnowNo, as well as all other baselines, with a higher success-to-help ratio across threshold values. Quantitatively, LAP+$a_{c}$ reduces the amount of help needed by over 70%, from 30.1% with KnowNo to 8.9% at a success rate of 60%. Similarly, prediction sets created by LAP are smaller than those from KnowNo. On the Human Help Rate curve, we find that LAP+$a_{c}$ has an Auc of $0.780$ compared to KnowNo with an AuC of $0.699$.
Of Prompt and Binary, both produce a singular success rate and help rate. Binary performed the best for this environment, getting close to the performance of LAP+$a_{c}$ at its success level of 77%. Prompt performs relatively poorly, with a success rate below 60%, illustrating the challenges present when attempting to retrieve the prediction sets directly from the LLM.
LAP with the perception-based A-Feasibility metric, LAP+$a_{perc}$ also under-performs LAP with the context-based score. We believe this is caused by MCQA options, $y^{i}$, becoming more difficult to separate relative to LAP+$a_{c}$, which introduces certainty. We believe that $a_{perc}$ may still be valuable in deployed scenarios without perfect perception. Similarly, LAP+$a_{pr}$ exhibited marginally lower performance than the context-based A-Feasibility metric. We believe this is also due to the context-based scores certainty helping its performance in this simple task.
Figure 5: Example of our real-world experimentation setup. The ClearPath TurtleBot must navigate towards the correct object, at which point a human moves the object into the basket before the robot continues towards a goal location. We chose this setup as no viable mobile manipulator was available to us.
IV-B Real World: Mobile Manipulator in a Kitchen
For this environment, we continue to use the task specifications laid out in KnowNo [3]. Shown in Figure 5, each scenario involves a mobile robot in front of a countertop in an office kitchen, next to a set of recycling/compost/landfill bins. The tasks include picking up objects from the countertop and possibly putting it into a bin, or somewhere else on the countertop. New ambiguities are added relative to the simulator environment, including some involving unsafe actions (e.g. task: “place the bowl in the microwave.”)The full outline of every possible ambiguity is included in the supplementary material. In our lab, no viable mobile manipulator exists, so we run our testing on a ClearPath TurtleBot. During the testing, we manually pick up the object selected by the agent and put it in a basket on the robot, simulating the use of the manipulator. We mark a task a success if the robot indicates the correct actions, moves to within 1 meter of the target object or destination, and orients itself towards the target.
Similar to the simulation environment, we find that LAP outperforms KnowNo and our other baselines. We show these results in Figure 4. Our prompt-based metric and context-based metric combined into a superscore, LAP+$a_{pr}\times a_{c}$ produced our best results, with a human help rate of only 42% relative to KnowNo’s 63% at a 70% planning success rate. These findings are corroborated with LAP having a lower average prediction set size than KnowNo, and an improved AuC of $0.72$ relative to KnowNo’s $0.67$. Qualitatively, we notice that LAP+$a_{pr}\times a_{c}$ improves performance on scenarios with a safety component, likely due to $a_{pr}$ scoring each option partially on its safety. We believe the performance of LAP+$a_{pr}\times a_{c}$ is due to each term contributing differently, with $a_{pr}$ measuring not just possibility, but also safety, for a given action while $a_{c}$ removing many hallucinated objects or destinations from the list of available options entirely. We can identify a similar trend for the Binary and Prompt baselines as in the simulation environmentwith the added complexity and ambiguities of the environment further hampering the Prompt baseline’s ability to produce an accurate prediction set directly.
Different from KnowNo, we ran additional experiments without perfect perception where we found the scene context through the VLM LLaVA. We then produced the perception-based A-Feasibility metric LAP+$a_{perc}$ using the objects listed in the LLaVA output. We only tested this on 20 scenarios due to the manual labor needed to verify LLaVA contexts and subsequent selected options versus the constrained KnowNo scenarios. We found that LAP+$a_{perc}$ was very useful to screen objects that were hallucinated by LLaVA which occurred in 4 of our scenarios. Otherwise, the context-based metric continued to perform better, with all advantaged by the use of the prompt-based metric.
Figure 6: Comparison of LAP+$a_{pr}\times a_{c}$ and KnowNo accross different LLMs. Note that LAP improves upon KnowNo with every model, and that performance generally increases with each model.
IV-C LLM Ablations
To further test LAP, we ablated our model choice, GPT-4 for both our method, LAP+$a_{pr}\times a_{c}$, and KnowNo in our real-world paradigm. Figure 6 shows these results. We show a large improvement over KnowNo across all models. The performance gap between LAP and KnowNo is wider for GPT-3.5-Turbo and GPT-4-Turbo than it is for GPT-4. We hit maximum performance on GPT-4-Turbo at half the human help rate of KnowNo. Additionally, we show that the performance of LAP improves with the model.
As an additional ablation, we also tested LAP using the ablated model for the MCQA generation and scoring phases, but only GPT-3.5-Turbo for finding $a_{pr}$. Performance was largely similar with help rate increasing by under 4% on average. We envision this as a cost-saving measure for users of LAP who want to minimize API or compute costs. These plots can be seen in our supplementary materials.
IV-D Why Not Just Fine-Tune?
In simulation, where we can generate a large set of possible tasks, we attempted to fine-tune gpt-3.5-turbo-1106, the best model available at the time, on each part of LAP. To test generalization, we separate out certain ambiguities of each type, as well as one entire type, for use exclusively in the testing set, which has a distribution identical to our dataset utilized above.
We find that overall help rates increase for the same model by on average 14%. This difference is largely due to impacted performance on ambiguities that were outside of the training data, showing a tendency to overfit seen by others including [3]. Similar issues occurred when fine-tuning on the option scoring component of LAP, decreasing help rates by less than 5% on average, but requiring help more frequently on the withheld ambiguities. This method still underperforms LAP on GPT-4 across most success rates.
In the real-world environment, with even more limited data, fine-tuning the MCQA proposal produced an overfitted model that could not generalize to the testing set of tasks.When finetuning on scoring, help rates decreased on average but with significantly worse performance in withheld categories. The fine-tuned models still underperform LAP on GPT-4. We can draw two separate conclusions from these experiments: (1) Generating a large and diverse enough dataset of scenarios in a robotic paradigm is exceptionally difficult, and (2) The newest, most powerful, models available outperform a fine-tuned older model for this task.
V Conclusions, Limitations, and Future Work
Robots must complete user-provided instructions accurately, with minimal instruction from the user. In this work we introduce LAP, a method of using an action feasibility score to significantly reduce the impact of LLM hallucinations and minimize the amount of human feedback needed when completing tasks. This includes a reduction in human help over prior art KnowNo [3] of 33% at a success rate of 70% in our real-world experiments.
Limitations and Future Work. The primary limitation of the methods we propose is perception. While providing ground-truth perception follows prior art, the uncertainty in perception found in a real-world environment is important to properly handle. An additional limitation is the limited diversity of the dataset we tested on, mainly that it only includes pick-and-place tasks. Future work on more diverse datasets would be valuable.
Appendix A Additional Experiment Details
Full Prompt for Prompt-Based A-Feasibility
{Header}
We: On the counter, there is a metal bowl, a plastic bowl, and a microwave.
We: pick up the metal bowl and put it in the microwave
We: Is this possible and safe given the provided knowledge of the scene?
You: False
We: On the counter, there is an orange, a bag of rice chips, and an apple.
We: pick up the orange
We: Is this possible and safe given the provided knowledge of the scene?
You: True
We: On the counter, there is
$\mathcal{O}$}.
We:
{$y^{i}$}
We: Is this possible and safe given the provided knowledge of the scene?
You:
MCQA Prompt. ADD HERE
Simulation setting. This information on the simulated task is directly from [3], the source of the task.
  • Environment: there are always three blocks and bowls of color red, yellow, and green with random locations on the table.
  • Goal: we use the following template: put, place, move a, one, a single of, two, a pair of, three, all, red, yellow, green block(s), bowl(s) on, to the left of, to the right of, to the front of, at the back of the red, green, yellow block(s), bowl(s). The scenario distribution is uniform over these possibilities.
  • Instruction: for the language instructions, we modify the goal (from the template) according to the ambiguity type. The scenario distribution is uniform over the listed ambiguous cases in each ambiguity type.
    • Attribute ambiguities: refer to the block as one of “cube”, “cuboid”, “box”, “square object”, to the bowl as one of “container”, “round object”, “receptacle”, or to either block or bowl as one of “object”, “item”, “thing” (“move the blue object in yellow bowl”); refer to “blue” as one of “cyan”, “navy”, to “green” as one of “greenish”, “grass-colored”, and to “yellow” as “orange” or “gold”. This setting is the least ambiguous one among the three ambiguity types.
    • Numeric ambiguities: refer to either two or three numerically with one of “a few”, “a couple of”, “some”, “a handful of” (“put some blocks in the green bowl”).
    • Spatial ambiguities: refer to any of the four possible directions with “near”, “close to”,“beside”, “next to”, refer to either left to right with “lateral to”, and refer to either front or behind with “along the line of sight”. This setting is the most ambiguous one among the three ambiguity types.
Outsize of [3]’s task description, we exchanged communications with them that provided further clarification for setting up this task. Our additional findings which should help with setting up this task are below.
  • For goals with numeric ambiguities, the ‘block’ is always the object to move, and all the ‘blocks’ are the same color to avoid multiple ambiguities.
  • The authors of [3] used different few shot prompts for each ambiguity, so that the examples provided matched the ambiguity accordingly.
We chose not to conform to the second point, instead using one few shot prompt for all ambiguities. We believe that this is more faithful to the task as the agent should have no knowledge of which ambiguity is being provided to it. We did test multiple prompts with varying examples and chose the one with the best results, although results did not change much when using reasonable examples of all ambiguity types. Our code, which will be release after acceptance, will show how we produced the full dataset listing from [3] and should help resolve any additional questions.
Hardware Mobile Manipulator setting. This information on the mobile manipulator task is directly from [3], the source of the task.
  • Environment: the full list of possible objects include: bottled water, bottled tea, orange soda, RedBull, Coke, Pepsi, Sprite, rice chips, jalapeno chips, kettle chips, multigrain chips, apple, orange, energy bar, clean sponge, dirty sponge, metal bowl, plastic bowl. Depending on the ambiguity listed below, there is three objects placed on the top of the counter (including randomly sampled distractors from the list). There is also a set of landfill, compost, and recycling bins, a microwave, and a portable stove.
  • Instruction: for convenience, we introduce the possible instructions first in different ambiguous scenarios; they each correspond to possible goals. Please refer to https://robot-help.github.io/prompts/mobile_tasks.txt for the full list. The possible instructions are a uniform distribution over different types: (1) single-label, e.g., ‘Bring me a Coke’ (unambiguous); (2) creative-single-label, e.g., ‘I want a healthy fruit to munch on.’ which means the apple (unambiguous); (3) multi-label, e.g., ‘Bring me a cola.’ and either Coke or Pepsi is acceptable; (4) creative-multi-label, e.g., ‘Bring me something with a kick.’ and either RedBull or jalapeno chips are acceptable; (5) spatially-ambiguous, e.g., ‘Put the Coke in the drawer’ or ‘Put the Coke near the fruit’ which under-specifies the drawer or fruit; (6) unsafe, e.g., ‘Can you dispose of the bottle drink? It should have expired.’ or ‘Place the bowl on the stove, please.’; (7) Winograd, e.g., ’There is a sponge and a bag of rice chips…I don’t want to use it for cleaning any more. Can you please dispose of it?” We use the GPT-4 model for generating the creative tasks.
  • Goal: the corresponding goal for the ambiguous instructions above. For example, the instruction is “Put the Coke in the drawer”, and the goal is uniform over the two possibilities: put the Coke in the top drawer, and put the Coke in the bottom drawer.
We faithfully implement [3]’s task description exactly as described, using their code provided here: https://github.com/google-research/google-research/tree/master/language_model_uncertainty. They specifically provide two files with their 300 tasks and corresponding prompts in the code. While this task is challenging and diverse, we believe that more can be done to expand the task to a much larger set of commands and situations. This is very limited in scope with very few object classes, and most commands boiling down to picking up and moving an object. Future work must be conducted to expand this task for more general future robots.
Appendix B Additional LLM Ablations
As described in our experimentation section, we tested LAP with one model (GPT-4, GPT-4-Turbo, and GPT-3.5-Turbo) for the MCQA generation and scoring, but GPT-3.5-Turbo for finding our prompt-based A-Feasibility score, $a_{pr}$. This is a possible cost saving measure for users of LAP who need to minimize their API or compute costs. This plot is found in Figure 7. We find that performance slightly decreases when using GPT-3.5-Turbo for our prompt-based A-Feasibility score. We suggest that users run some testing on their own use case to determine the right trade off.
Figure 7: Comparison of LAP+$a_{pr}\times a_{c}$ and KnowNo accross different LLMs. Note that LAP improves upon KnowNo with every model, and that performance overall generally increases with each model. GPT-4-Turbo exhibited poor performance in the MCQA generation phase relative to GPT-4, frequently asking for clarifications instead of producing a set of candidate options, capping the possible success rate.